Solve Problems Using Linear Systems
9. Milk and cream contain different percents of butterflat. How much 3% milk needs to be mixed with how much 15% cream to give 20 L of 6% cream?
(answers are: 15 L of 3% milk, 5 L of 15% cream, but i dont know how to solve it?)

Answer by stanbon(75887)

Milk and cream contain different percents of butterfat. How much 3% milk needs to be mixed with how much 15% cream to give 20 L of 6% cream?
---------------------
Equation:
fat + fat = fat
0.03x + 0.15(20-x) = 0.06*20
3x + 15*20 - 15x = 6*20
-12x = -9*20
x = (3/4)*20
x = 15 L (amt. of 3% milk in the mixture)
20-x = 5 L (amt. of 15% cream in the mixture)
==================================================
Cheers,
Stan H.